OpenStudy (anonymous):

Determine the slope of the line that passes through (-3,-4) and (1,2). Show your work.

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = (2-(-4))/(1-(-3)) m = (2+4)/(1+3) I'll let you finish

jimthompson5910 (jim_thompson5910):

Don't forget to reduce any/all fractions

OpenStudy (anonymous):

m= 1/2?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

close, but not quite

jimthompson5910 (jim_thompson5910):

m = (2+4)/(1+3) m = 6/4 m = ??? .... reduce the last fraction above

OpenStudy (anonymous):

1.5?

jimthompson5910 (jim_thompson5910):

keep it as a fraction

jimthompson5910 (jim_thompson5910):

but reduce as much as possible

OpenStudy (anonymous):

1/2

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

6 and 4 have 2 in common as a factor 6/2 = 3 4/2 = 2 So 6/4 reduces to 3/2